# A Survey on Autonomous Driving Datasets: Data Statistic, Annotation, and Outlook

### v1

@@ VII-B Chronological Trends in Perception Datasets
In Fig. 10, we introduce a chronological overview of perception datasets with the top 50 impact scores from 2007 to 2023 (until the writing of this paper). The datasets are color-coded according to their sourcing domain, and synthetic datasets are marked with an external red outline, clearly illustrating the progress toward the diverse data collection strategy. A noticeable trend shows the increase in the number and variety of datasets over the years, indicating the requirement for high-quality datasets with the growing advancements in the field of autonomous driving.
In general, most of the datasets provide a perception perspective from the sensors equipped on the ego vehicle (onboard) because of the importance of the capability of the autonomous vehicle to efficiently and precisely precept the surroundings. On the other hand, due to the high-cost real-world data, some researchers propose high-influence synthetic datasets like VirtualKITTI [3] (2016) to alleviate the reliance on real data. Facilitated by the effectiveness of simulators, there are many novel synthetic datasets [7] [10] published in recent years. In the timeline, V2X datasets like DAIR-V2X [27] (2021) also exhibit a trend toward cooperative driving systems. Furthermore, because of the non-occlusion perspective provided by UAV, drone-based datasets, such as UAVDT [87] published in 2018, take a crucial position in advancing perception systems.
@@ VII-C Data Distribution
Beyond the number of objects in a scene, the object distribution based on the distance to the ego vehicle is another essential point for revealing a dataset’s variety and significant differences, illustrated in Fig. 15. The Waymo dataset demonstrates numerous labeled objects in near-field to mid-field scenarios. In contrast, Argoverse 2 and ZOD show a wider detection range, with some frames even including bounding boxes out of 200 meters. The curve of the nuScenes means it is particularly rich in objects in a shorter range, which is typical for urban driving scenarios. Nevertheless, as the distance increases, the nuScenes dataset quickly tappers off for the number of objects with annotations. The ONCE dataset covers a more even distribution of objects across distances, while the KITTI dataset focuses more on close-range detection.
@@ VIII Discussion and Future Works
NuScenes-QA [297] 2023 34K scenes, 460K QA pairs $\checkmark$ onboard visual question answering real
@@
Driving LLMs [298] 2023 10K driving situations, 160K QA pairs $\times$ drone visual question answering synthetic
DRAMA [299] 2023 17,785 scenarios $\checkmark$ onboard reasoning, visual question answering real
@@
NuScenes-MQA [303] 2023 34K scenarios, 1.4M QA pairs $\times$ onboard visual question answering real
@@
TABLE VII: VLM autonomous driving dataset. OT: object tracking, MOT: multi-object tracking, QA: question-answering
@@
In this paper, we mainly focus on analyzing the existing datasets, which usually include rich visual data, and aim at tasks in the modular-based pipeline. However, with the rapid technological development, especially the excellent performance of the Large Language Models, many novel trends of the next-generation autonomous driving datasets have occurred while proposing new challenges and requirements.
@@
End-to-End Driving Datasets. Compared to the modular-designed autonomous driving pipeline, the end-to-end architecture simplifies the overall design process and reduces integration complexities. The success of the UniAD [307] verifies the potential ability of end-to-end models. However, the number of datasets for end-to-end AD is limited [167, 113]. Therefore, introducing datasets focusing on end-to-end driving is crucial for advancing autonomous vehicles. On the other hand, implementing an automatic labeling pipeline in a data engine can significantly facilitate the development of end-to-end driving frameworks and data [67].
@@
Introduce Language into AD Datasets. The vision language models (VLMs) have recently achieved impressive advancement in many fields. Its inherent advantage in providing language information to vision tasks makes autonomous driving systems more explainable and reliable. [308] highlighted the prominent role of Multimodal Large Language Models in various AD tasks, such as perception [290, 296], motion planning [301], and motion control [111]. The autonomous driving datasets including language labels are shown in Tab. VII. Overall, incorporating language into AD datasets is a trend of the future development of AD datasets.
@@
Data Generation via VLMs. As mentioned by [309], the powerful capability of VLMs can be used to generate autonomous driving data. For example, DriveGAN [310] generated high-quality AD data by disentangling different components without supervision. Additionally, due to the capability of world models for comprehending driving environments, some works [311, 312, 313] explored world models to generate high-quality driving videos. DriveDreamer[312] as a pioneering work derived from real-world scenarios, addressing the limitation of the gaming environments or simulated settings.
@@
Domain Adaptation. Domain adaptation is a critical challenge in developing autonomous vehicles [314], referring to the ability of a model trained on one dataset (the source domain) to perform stable on another dataset (the target domain). This challenge manifests in multiple aspects, such as diversity in environmental conditions [315], sensor settings [316], or synthetic-to-real transform [317].
@@ IX Conclusion
In this paper, we exhaustively and systematically reviewed and analyzed more than 200 existing autonomous driving datasets. We started with the sensor types and modalities, sensing domains, and tasks relevant to autonomous driving datasets. We introduced a novel evaluation metric called impact score to validate the influence and importance of perception datasets. Afterward, we demonstrated several high-impact datasets involving perception, prediction, planning, control, and end-to-end autonomous driving. Additionally, we explained the annotation methodology for autonomous driving datasets and investigated the factors affecting annotation quality.
Moreover, we delineated the chronological and geographical distribution of the collected datasets, providing a comprehensive perspective for understanding the current development of autonomous driving datasets. Meanwhile, we studied the data distribution of several datasets, offering a specific viewpoint for comprehending the variances across different datasets. In the end, we discussed the development and trend of the next generation of autonomous driving datasets.

### v2

@@ VII-B Chronological Trends in Perception Datasets
In Fig. 10, we introduce a chronological overview of perception datasets with the top 50 impact scores from 2009 to 2024 (until the writing of this paper). The datasets are color-coded according to their sourcing domain, and synthetic datasets are marked with an external red outline, clearly illustrating the progress toward the diverse data collection strategy. A noticeable trend shows the increase in the number and variety of datasets over the years, indicating the requirement for high-quality datasets with the growing advancements in the field of autonomous driving.
In general, most of the datasets provide a perception perspective from the sensors equipped on the ego vehicle (onboard) because of the importance of the capability of the autonomous vehicle to efficiently and precisely precept the surroundings. On the other hand, due to the high-cost real-world data, some researchers propose high-influence synthetic datasets like VirtualKITTI [3] (2016) to alleviate the reliance on real data. Facilitated by the effectiveness of simulators, there are many novel synthetic datasets [7] [10] published in recent years. In the timeline, V2X datasets like DAIR-V2X [27] and TUMTraf ([84, 85, 86, 60]) also exhibit a trend toward cooperative driving systems. Furthermore, because of the non-occlusion perspective provided by UAV, drone-based datasets, such as UAVDT [94] published in 2018, take a crucial position in advancing perception systems.
@@ VII-C Data Distribution
Beyond the number of objects in a scene, the object distribution based on the distance to the ego vehicle is another essential point for revealing a dataset’s variety and significant differences, illustrated in Fig. 15. The Waymo dataset demonstrates numerous labeled objects in near-field to mid-field scenarios. In contrast, Argoverse 2 and ZOD show a wider detection range, with some frames even including bounding boxes out of 200 meters. The curve of the nuScenes means it is particularly rich in objects in a shorter range, which is typical for urban driving scenarios. Nevertheless, as the distance increases, the nuScenes dataset quickly tappers off for the number of objects with annotations. The ONCE dataset covers a more even distribution of objects across distances, while the KITTI dataset focuses more on close-range objects.
VII-D Influence of Adversarial Environmental Conditions.
TABLE IV: 3D object detection performance under adversarial environmental conditions on the nuScenes [22] dataset. We report mAP (%) as the evaluation metric. ’L’ and ’C’ refer to LiDAR and cameras, respectively. The term ’Normal’ indicates that the model was tested across the entire validation set, encompassing all environmental conditions.
Method Modality Adversarial Conditions mAP $\uparrow$
VoxelNext [170] L normal 60.5
nighttime 32.4
rainy 58.7
UVTR [171] L normal 60.8
nighttime 33.9
rainy 56.5
Transfusion [172] L+C normal 68.9
nighttime 37.5
rainy 65.8
We further study the influence of adversarial environmental conditions (low lightning and rain) on the performance of 3D object detectors in the autonomous driving system. The experiment results are reported in Tab. IV. We utilize the nuScenes [22] dataset and choose three state-of-the-art methods, VoxelNext [170], UTVR [171], and Transfusion [172]. For a fair comparison, we directly utilize the pre-trained checkpoints provided in the repository of each approach. The rainy and nighttime data are also manually collected from the nuScenes validation set. All three methods demonstrate similar trends under different environmental conditions. Specifically, the detection accuracy declines significantly under low illumination conditions compared with evaluations of the whole validation set. Besides, all models exhibit slight decreases under the non-heavy rainy situations recorded by nuScenes. The reliability of detectors in the real world could be much worse once the rain is heavy. Therefore, taking image restoration into account is a promising way for camera or sensor fusion-based approaches to overcome these challenges [173, 174]. In conclusion, increasing the amount of data recorded on various types and intensities of weather conditions is critical for training a robust and reliable autonomous driving system.
@@ VIII Discussion and Future Works
TABLE V: VLM autonomous driving dataset. OT: object tracking, MOT: multi-object tracking, QA: question-answering, DM: decision-making, IP: intention prediction, VR: visual reasoning, SR: spatial reasoning
@@
NuScenes-QA [183] 2023 1K scenarios, 460K QA pairs $\checkmark$ onboard visual QA real
@@
Driving LLMs [184] 2023 10K driving situations, 160K QA pairs $\times$ drone visual QA synthetic
DRAMA [185] 2023 17,785 scenarios $\checkmark$ onboard reasoning, visual QA real
@@
NuScenes-MQA [189] 2023 34K scenarios, 1.4M QA pairs $\times$ onboard visual QA real
@@
Talk2BEV [193] 2023 1K BEV scenarios, 20K QA pairs $\checkmark$ onboard DM, IP, VR, SR real
NuScenes-MQA [194] 2024 1.4M QA pairs, 34,149 scenarios $\checkmark$ onboard visual QA real
VLAAD [195] 2024 10,379 scenarios $\checkmark$ onboard visual QA, reasoning real
@@
With rapid technological development, powerful computation resources, and excellent artificial intelligent algorithms, many novel trends in next-generation autonomous driving datasets have occurred while proposing new challenges and requirements.
@@
End-to-End Driving Datasets. Compared to the modular-designed autonomous driving pipeline, the end-to-end architecture simplifies the overall design process and reduces integration complexities. The success of UniAD [196] verifies the potential ability of end-to-end models. However, the number of datasets for end-to-end AD is limited [129]. Therefore, introducing datasets focusing on end-to-end driving is crucial for advancing autonomous vehicles. On the other hand, implementing an automatic labeling pipeline in a data engine can significantly facilitate the development of end-to-end driving frameworks and data [74].
Potential Applications of AD Datasets Future autonomous driving datasets should provide extensive real-world environmental and traffic data, supporting a wide range of applications beyond the ego-vehicle and basic vehicle-infrastructure cooperation. For instance, the interaction data between autonomous vehicles and intelligent infrastructure components can guide the advancement of Internet-of-Thing (IoT)-enabled devices like smart traffic signals. Moreover, the detailed insights into traffic patterns, congestion, and vehicle behavior across different times and conditions can facilitate urban planning, optimize the traffic flow, and enhance overall traffic management strategies.
@@
Introduce Language into AD Datasets. Vision language models (VLMs) have recently achieved impressive advancement in many fields. Its inherent advantage in providing language information to vision tasks makes autonomous driving systems more explainable and reliable. The survey [197] highlights the prominent role of Multimodal Large Language Models in various AD tasks, such as perception [176, 182], motion planning [187], and motion control [193]. Autonomous driving datasets including language labels are shown in Tab. V. Overall, incorporating language into AD datasets is a trend of the future development of AD datasets.
@@
Data Generation via VLMs. The powerful capability of VLMs can be used to generate data [198, 199, 200, 201]. For example, DriveGAN [202] generated high-quality AD data by disentangling different components without supervision. Additionally, due to the capability of world models for comprehending driving environments, some works [203, 204, 205] explored world models to generate high-quality driving videos. DriveDreamer [204] as a pioneering work derived from real-world scenarios, addressing the limitation of gaming environments or simulated settings. The most recent advancements in text-video generation, exemplified by Sora [206], have opened a new avenue for generating autonomous driving data. With a brief description, Sora can generate high-quality synthetic data that is very close to real scenes. This capability significantly enhances dataset augmentation, especially by extending the data available for rare events like traffic accidents. The increased data availability provided by VLMs is poised to enhance the training and evaluation of autonomous driving systems, potentially improving their safety and reliability.
@@
Domain Adaptation. Domain adaptation is a critical challenge in developing autonomous vehicles [207], referring to the ability of a model trained on one dataset (the source domain) to perform stable on another dataset (the target domain). This challenge manifests in multiple aspects, such as diversity in driving conditions [208], sensor settings [209], or synthetic-to-real transform [210]. Therefore, a trend of developing the next-generation datasets is incorporating data from heterogeneous sources. First of all, datasets should not only cover a wide range of environmental conditions (various weather conditions and day-night illumination) but also include diverse geographic locations. Second, combining data from various sensor types is also pivotal to overcoming the domain adaptation issue. Another solution is taking into account blending high-quality synthetic data and real-world data in a balanced way to improve generalization.
Uncertainty Issues in Autonomous Driving. Usually, uncertainty is modeled in a probabilistic way in the field of machine learning. The uncertainty can be referred to two types: a) aleatoric uncertainty, where the uncertainty like noise stems from the data, and b) epistemic uncertainty refers to uncertainty caused by unawareness about the best model [211]. For autonomous driving, one of the major reasons resulting in uncertainty is the incompleteness of training data [212]. The insufficient training data can not wholly represent the driving environment, causing autonomous vehicles to problematically deal with rare cases during operation. Therefore, improving the diversity of the dataset for training reliable autonomous driving systems is crucial. Moreover, datasets include rare events and edge cases, allowing models to better understand and quantify the uncertainty associated with unexpected situations and safely handle them.
Standardization of Data Creation. Addressing the standardization of data creation is critical aspect for developing new datasets, as it directly impacts the accuracy, efficiency, and reliability of autonomous vehicle models. In general, data standardization refers to the attributes of the data, terminology and structure of the dataset, and data storage [213]. For the data attributes, unifying data formats across different sensor types and source domains is valuable to facilitate integrated processing and analysis. Developing comprehensive guidelines for dataset labeling is crucial to guarantee consistent, high-quality annotations, enhancing model training across diverse datasets and boosting performance reliability. Moreover, establishing standardized data storage and access protocols is vital, enabling seamless dataset sharing and integration across multiple sources, and fostering collaboration within the autonomous driving research community.
Data Privacy. The advancement of autonomous vehicles require a lot of data to guarantee the driving safety. However, the more the available data, the greater the concern that private data will be abused. For the earlier autonomous driving datasets like KITTI [13], there is no anonymization progress for the recorded data, especially for images, which leads to a potential risk for leak private information. With the introduction and refinement of relevant regulations in various countries, data anonymization has been widely adopted in recent datasets [84, 22, 23]. Nevertheless, the information provided by a large dataset is more than individual biographic features or vehicle plates. Professional institutes can indirectly extract external information by analyzing the existing information in the data. For instance, by analyzing the types of vehicles and the attire of pedestrians in the dataset, one can infer information about the infrastructure, basic construction, and other characteristics of the surrounding area in which the data were collected.
Open Data Ecosystems. The primary objective of open data ecosystems (ODE) for autonomous driving is to foster innovation, enhance transparency, and facilitate collaboration across governments, companies, and research communities. This is achieved through the free exchange of datasets, breaking down the traditional barriers that have restricted access to corporations and research institutions. By doing so, ODE empowers a wider range of innovators to participate in autonomous driving development, thereby boosting a more diverse and inclusive innovation ecosystem. Additionally, ODE establishes dynamic feedback loops where users can report issues, propose improvements, and contribute to enhancing datasets. Nevertheless, unrestricted access to data raises strong concerns about data security and privacy. Addressing these concerns necessitates carefully crafting and continuously refining relevant legal frameworks to safeguard sensitive information while prompting the growth of ODE. The balance is pivotal for maintaining the integrity of open data ecosystems and ensuring their sustainable development in autonomous driving.
@@ IX Conclusion
In this paper, we exhaustively and systematically reviewed 265 existing autonomous driving datasets. We started with the sensor types and modalities, sensing domains, and tasks relevant to autonomous driving. We introduced a novel evaluation metric called impact score to validate the influence and importance of perception datasets. Our analysis covered the attributes and significance of high-impact datasets across perception, prediction, planning, control, and end-to-end driving tasks. Additionally, we investigated the annotation methodologies and the factors affecting annotation quality. We also analyzed datasets from chronological and geographical perspectives to understand current trends and conducted experiments demonstrating the crucial need for data diversity under adversarial conditions. With the study on data distribution, we offered a specific viewpoint for comprehending the variances across different datasets. Our findings underscore the essential role of diverse, high-quality datasets in shaping the future of autonomous driving. Looking forward, we highlighted key challenges and future directions for autonomous driving datasets, including the adoption of VLM, domain adaptation, uncertainty challenges, standardization, data privacy, and open data ecosystems. These areas represent promising avenues for future research and are pivotal for advancing autonomous driving technology, setting a clear roadmap for further innovation in this rapidly evolving field.
Supplementary Tables of AD Datasets
TABLE VI: Part I of Autonomous Driving Datasets. SS: semantic segmentation, OD: object detection, OT: object tracking, MOT: multi-object tracking
Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
TUD-Brussels [215] 2009 1,600 frames $\times$ onboard (3D) OD real
TME Motorway [221] 2012 30K frames $\checkmark$ onboard OD, OT real
MSLU [222] 2013 36.8 km distances $\checkmark$ onboard SLAM real
NYC3DCars [225] 2013 2K images $\times$ onboard OD real
AMUSE [226] 2013 117,440 frames $\checkmark$ onboard SLAM real
Paris-rue-Madame [229] 2014 643 objects $\times$ onboard OD, SS real
Ua-detrac [232] 2015 140K frames $\checkmark$ V2X OD, MOT real
NCLT [233] 2015 34.9 hours $\checkmark$ onboard odometry real
SAP [236] 2016 19K frames $\times$ drone OD, OT real
TRoM [241] 2017 712 frames $\times$ onboard road marking detection real
NEXET [242] 2017 91,190 frames $\times$ onboard OD real
DIML [243] 2017 470 videos $\checkmark$ onboard lane detection real
Complex Urban [245] 2017 around 190km paths onboard SLAM real
Syncapes [249] 2018 25K images $\times$ onboard OD, SS synthetic
NVSEC [250] 2018 around 28 km distances $\checkmark$ others SLAM real
CADP [252] 2018 1,416 scenes $\times$ V2X traffic accident analysis real
nighttime drive [255] 2018 35K $\times$ onboard SS real
VEIS [160] 2018 61,305 frames $\times$ onboard OD, SS synthetic
EuroCity Persons [258] 2018 47,300 frames $\times$ onboard OD real
RANUS [259] 2018 4K frames $\times$ onboard SS, scene understanding real
D2-City [261] 2019 700K annotated frames $\checkmark$ onboard OD, MOT real
Mcity [263] 2019 1,7500 frames $\checkmark$ onboard SS real
PreSIL [8] 2019 50K frames $\checkmark$ onboard OD, 3D SS synthetic
H3D [265] 2019 27,721 frames $\checkmark$ onboard (3D) OD, MOT real
LLAMAS [266] 2019 100K frames $\checkmark$ onboard lane segmentation real
MIT DriveSeg [267] 2019 5K frames $\times$ onboard SS real
Astyx [268] 2019 500 radar frames $\times$ onboard 3D OD real
UNDD [269] 2019 7.2K frames $\checkmark$ onboard SS real
Boxy [270] 2019 200K frames $\checkmark$ onboard OD real
RUGD [271] 2019 37K frames $\checkmark$ others SS real
Brno Urban [276] 2019 375.7 km $\checkmark$ onboard recognition real
VLMV [279] 2020 306K frames $\checkmark$ V2X lane merge real
Cirrus [281] 2020 6,285 frames $\checkmark$ onboard (3D) OD real
A*3D [283] 2020 39,179 point cloud frames $\times$ onboard (3D) OD real
Toronto-3D [284] 2020 4 scenarios $\times$ onboard 3D SS real
CADC [168] 2020 56K $\times$ onboard (3D) OD, OT real
IDDA [287] 2020 1M frames $\times$ onboard segmentation synthetic
TABLE VII: Part II of Autonomous Driving Datasets
Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
CARRADA [288] 2020 7,193 radar frames $\checkmark$ onboard SS real
NightCity [290] 2020 4,297 frames $\times$ onboard nighttime SS real
DDAD [291] 2020 21,200 frames $\times$ onboard depth estimation real
CPIS [92] 2020 10K frames $\times$ V2X cooperative 3D OD synthetic
EU LTD [294] 2020 around 37 hours $\checkmark$ onboard odometry real
CCD [296] 2020 4.5K videos $\checkmark$ onboard accident prediction real
DSEC [47] 2021 53 sequences (3193 seconds in total) $\checkmark$ onboard dynamic perception real
4Seasons [300] 2021 350km recordings $\checkmark$ onboard SLAM real
PVDN [301] 2021 59,746 frames $\checkmark$ onboard nighttime OD, OT real
DRIV100 [302] 2021 100K frames $\times$ onboard domain adaptation SS real
NEOLIX [303] 2021 30K frames $\checkmark$ onboard 3D OD, OT real
IPS3000+ [165] 2021 14,198 frames $\checkmark$ V2X 3D OD real
DurLAR [305] 2021 100K frames $\checkmark$ onboard depth estimation real
Multifog KITTI [309] 2021 15K frames $\times$ onboard 3D OD synthetic
Comap [310] 2021 4,391 frames $\checkmark$ V2X 3D OD synthetic
WIBAM [312] 2021 33,092 frames $\checkmark$ V2X 3D OD real
RaidaR [314] 2021 58,542 rainy street scenes $\times$ onboard SS real
ROAD [317] 2021 122K frames $\checkmark$ onboard OD, SS real
PandaSet [166] 2021 8,240 frames $\checkmark$ onboard (3D) OD, SS, OT real
SODA10M [318] 2021 20K labeled images $\checkmark$ onboard OD real
PixSet [319] 2021 29K point cloud frames $\times$ onboard (3D) OD real
VIL-100 [321] 2021 10K frames $\times$ onboard lane detection real
WADS [323] 2021 1K point cloud frames $\checkmark$ onboard SS real
CrashD [327] 2022 15,340 scenes $\times$ onboard 3D OD synthetic
RoadSaW [331] 2022 720K frames $\checkmark$ onboard Road surface and wetness estimation real
ASAP [333] 2022 1.2M images $\checkmark$ onboard online 3D OD real
Amodal Cityscapes [334] 2022 5K frames $\times$ onboard amodal SS real
TJ4RadSet [336] 2022 7,757 frames $\checkmark$ onboard OD, OT real
CODA [337] 2022 1,500 frames $\times$ onboard corner case detection real
AUTOCASTSIM [340] 2022 52K frames $\checkmark$ V2X (3D) OD, OT, SS real
K-Lane [342] 2022 15,382 frames $\times$ onboard lane detection real
Robo3D [347] 2023 476K frames $\checkmark$ onboard (3D) OD, 3D SS real
K-Radar [348] 2023 35K radar frames $\times$ onboard 3D OD, OT real
aiMotive [349] 2023 26,583 frames $\checkmark$ onboard 3D OD, MOT real
RoadSC [354] 2023 90,759 images $\times$ onboard road snow coverage estimation real
V2X-Real [355] 2024 171K images, 33K LiDAR point clouds - V2X 3D OD real
RCooper [356] 2024 50K images, 30K LiDAR point clouds $\checkmark$ V2X 3D OD, OT real